# General-Reasoner: Advancing LLM Reasoning Across All Domains

### v4

@@ 1 Introduction
Recent advancements in large language models (LLMs) have demonstrated strong potential in many expert-level tasks by following user instructions. One of the most significant breakthroughs in recent work is the use of reinforcement learning (RL), such as PPO and GRPO, to improve the reasoning capabilities of LLMs. A particularly noteworthy development is the “Zero” reinforcement learning setting proposed by DeepSeek-R1-Zero [1], which shows that training a base LLM directly via RL can unlock powerful reasoning capabilities without relying on a supervised fine-tuning step. RL for LLM reasoning has inspired multiple open-source efforts, such as SimpleRL [2], DAPO [3], and DeepScaleR [4, 5], which leverage variants of GRPO [6] to further enhance reasoning performance—either in the zero RL setting or via continuous reinforcement learning on models with supervised fine-tuning or distillation.
However, existing methods primarily focus on training and evaluation in mathematical reasoning or coding tasks. This narrow focus can be attributed to two main reasons: (1) Data Abundance: it is easier to harvest large-scale mathematical data from the Internet due to the abundance of international math and coding competitions and exams; (2) Answer Verification: mathematical domains allow for easy and reliable answer verification using straightforward rule-based methods (e.g., exact numeric matching, symbolic equation comparison), which provide accurate reward signals. However, this reliance on rule-based verifiers restricts the generalizability of the resulting models. Real-world reasoning often spans multiple disciplines—such as science, finance, and the humanities—and involves complex, long-tailed answer representations that cannot be reliably verified through heuristic rules. In fact, training solely on mathematical or coding data does not guarantee generalization to other reasoning domains. For example, although S1/S1.1 [7] significantly improves mathematical scores, it degrades performance on MMLU-Pro [8] by 4–6%.
@@
  • All-Domain Answer Verifier (General-Verifier): In addition to scaling data, existing rule-based verifiers struggle with the diverse answer types commonly found in broader domains—including short string answers, LaTeX expressions, and other structured forms—which limits the effectiveness of reasoning training beyond mathematics. We introduce a compact generative verifier model (1.5B parameters) explicitly trained to verify short answers in a context-aware, chain-of-thought manner, thereby providing robust and reliable reward signals for RL training. This approach improves the flexibility and scalability of RL training across diverse reasoning tasks.
@@ 2.1 Large Language Models for Reasoning
Advancements in large language models (LLMs) have demonstrated substantial potential in effectively performing diverse tasks by following user instructions. However, expert-level tasks—such as solving complex mathematical problems [12] or addressing practical STEM challenges [8]-require robust reasoning capabilities. Recent work on chain-of-thought reasoning has shown that LLMs can significantly enhance their performance by explicitly decoding their thought processes during inference [17]. Subsequent studies have further investigated this method, termed “test-time scaling,” to better understand and enhance the reasoning abilities of LLMs.
Following this direction, commercial models such as OpenAI’s O1 series have exhibited impressive performance on reasoning-intensive evaluations [18], including mathematical Olympiad problems. Meanwhile, recent open-source models—such as Qwen [19], QWQ [20], and Deepseek R1 [1]—have also achieved competitive results, narrowing the performance gap with state-of-the-art commercial counterparts. This progress opens new avenues for academia to explore advanced optimization strategies aimed at further improving the reasoning capabilities of LLMs [7, 3, 21].
@@ 4.4 Main Results
Stronger results are observed when initializing General-Reasoner with Qwen3 backbones. For example, General-Reasoner-4B surpasses Qwen2.5-7B after Zero RL, reaching 62.8% on MMLU-Pro versus 58.9%. This demonstrates the efficiency and transferability of our training method across model scales. The best performing version is General-Reasoner-Qw3-14B, which achieves 1% on GPQA and 54.4% on TheoremQA, matching the performance of commercial models like GPT-4o, which scores 50.0% and 43.6% respectively, despite relying solely on Zero RL. Compared to Qwen3-14B-Instruct (non-think), which undergoes post-training via distillation from a much bigger teacher model, our model maintains an advantage on many benchmarks. Notably, even in non-think mode, the Qwen3-Instruct model continues to generate CoT outputs. While some performance gaps remain relative to closed-source or closed-data systems, our results underscore the promise of Zero reinforcement learning when combined with domain-diverse reasoning data and a compact, model-based verifier. Although a performance gap remains on some benchmarks compared to closed-source or closed-data models, our results underscore the promise of Zero reinforcement learning when combined with domain-diverse reasoning data and a compact generative verifier.

### v5

@@ 1 Introduction
Recent advancements in large language models (LLMs) have demonstrated strong potential in many expert-level tasks by following user instructions. One of the most significant breakthroughs in recent work is the use of reinforcement learning (RL), such as PPO and GRPO, to improve the reasoning capabilities of LLMs. A particularly noteworthy development is the “Zero” reinforcement learning setting proposed by DeepSeek-R1-Zero [1], which shows that training a base LLM directly via RL can unlock powerful reasoning capabilities without relying on a supervised fine-tuning step. RL for LLM reasoning has inspired multiple open-source efforts, such as SimpleRL [2], DAPO [3], and DeepScaleR [4, 5], which leverage variants of GRPO [6] to further enhance reasoning performance, either in the zero RL setting or via continuous reinforcement learning on models with supervised fine-tuning or distillation.
However, existing methods primarily focus on training and evaluation in mathematical reasoning or coding tasks. This narrow focus can be attributed to two main reasons: (1) Data Abundance: it is easier to harvest large-scale mathematical data from the Internet due to the abundance of international math and coding competitions and exams; (2) Answer Verification: mathematical domains allow for easy and reliable answer verification using straightforward rule-based methods (e.g., exact numeric matching, symbolic equation comparison), which provide accurate reward signals. However, this reliance on rule-based verifiers restricts the generalizability of the resulting models. Real-world reasoning often spans multiple disciplines, such as science, finance, and the humanities, and involves complex, long-tailed answer representations that cannot be reliably verified through heuristic rules. In fact, training solely on mathematical or coding data does not guarantee generalization to other reasoning domains. For example, although S1/S1.1 [7] significantly improves mathematical scores, it degrades performance on MMLU-Pro [8] by 4–6%.
@@
  • All-Domain Answer Verifier (General-Verifier): In addition to scaling data, existing rule-based verifiers struggle with the diverse answer types commonly found in broader domains, including short string answers, LaTeX expressions, and other structured forms, limiting the effectiveness of reasoning training beyond mathematics. We introduce a compact generative verifier model (1.5B parameters) explicitly trained to verify short answers in a context-aware, chain-of-thought manner, thereby providing robust and reliable reward signals for RL training. This approach improves the flexibility and scalability of RL training across diverse reasoning tasks.
@@ 2.1 Large Language Models for Reasoning
Advancements in large language models (LLMs) have demonstrated substantial potential in effectively performing diverse tasks by following user instructions. However, expert-level tasks, such as solving complex mathematical problems [12] or addressing practical STEM challenges [8], require robust reasoning capabilities. Recent work on chain-of-thought reasoning has shown that LLMs can significantly enhance their performance by explicitly decoding their thought processes during inference [17]. Subsequent studies further investigate the scaling of reasoning process, termed “test-time scaling,” to better enhance the reasoning capability of LLMs [18, 19].
Following this direction, commercial models such as OpenAI’s O1 series have exhibited impressive performance on reasoning-intensive evaluations [20], including mathematical Olympiad problems. Meanwhile, recent open-source models, such as Qwen [21], QWQ [22], and Deepseek R1 [1], have also achieved competitive results, narrowing the performance gap with state-of-the-art commercial counterparts. This progress opens new avenues for academia to explore advanced optimization strategies aimed at further improving the reasoning capabilities of LLMs [7, 3, 23].
@@ 4.4 Main Results
Stronger results are observed when initializing General-Reasoner with Qwen3 backbones. For example, General-Reasoner-4B surpasses Qwen2.5-7B after Zero RL, reaching 62.8% on MMLU-Pro versus 58.9%. This demonstrates the efficiency and transferability of our training method across model scales. The best performing version is General-Reasoner-Qw3-14B, which achieves 56.1% on GPQA and 54.4% on TheoremQA, matching the performance of commercial models like GPT-4o, which scores 50.0% and 43.6% respectively, despite relying solely on Zero RL. Compared to Qwen3-14B-Instruct (non-think), which undergoes post-training via distillation from a much bigger teacher model, our model maintains an advantage on many benchmarks. Notably, even in non-think mode, the Qwen3-Instruct model continues to generate CoT outputs. While some performance gaps remain relative to closed-source or closed-data systems, our results underscore the promise of Zero reinforcement learning when combined with domain-diverse reasoning data and a compact, model-based verifier. Although a performance gap remains on some benchmarks compared to closed-source or closed-data models, our results underscore the promise of Zero reinforcement learning when combined with domain-diverse reasoning data and a compact generative verifier.